You and 11 coworkers just won $18 million ($1,500,000.00 each) from the state lottery. Assuming you each receive your share over 18 years and that the state lottery earns a 3 percent return on its funds, whatis the present value of your prize before taxes r you recu uest the up-front cash option? Click on the table icon to view the annuity table The present value of your prize before taxes if you request the up-front cash option is S Round your answer to the nearest dollar)
Answer #1

Each worker has won $ 1500000. Now the prize is paid out spread over a period of 18 years and the lottery earns 3 % on the fund.

Assuming the total prize is divided into equal sized payouts, each payout = 1500000 / 18 = $ 250000 / 3

Present Value of the Equal Sized Payouts = (250000/3) x (1/0.03) x [1-{1/(1.03)^(18)}] = $ 1146126.09

Answer: $ 1146126.09

NOTE: Total PV of Payouts of all winners = 12 x 1146126.09 = $ 13753513.08 ~ $ 13.753 million